# OxyGent: Making Multi-Agent Systems Modular, Observable, and Evolvable via Oxy Abstraction

### 28 Apr 2026 v1

@@ Abstract
Deploying production-ready multi-agent systems (MAS) in complex industrial environments remains challenging due to limitations in scalability, observability, and autonomous evolution. We present OxyGent, an open-source framework that enables modular, observable, and evolvable MAS via a unified Oxy abstraction, in which agents, tools, LLMs, and reasoning flows are encapsulated as pluggable atomic components. This Lego-like assembly paradigm supports scalable system composition and non-intrusive monitoring. To enhance observability, OxyGent introduces permission-driven dynamic planning that replaces rigid workflows with execution graphs generated at runtime, which provide adaptive visualizations. To support continuous evolution, the framework integrates OxyBank, an AI asset management platform that supports automated data backflow, annotation, and joint evolution. Empirical evaluations and real-world case studies show that OxyGent provides a robust and scalable foundation for MAS. OxyGent is publicly available at https://oxygent.jd.com/.
@@ 1 Introduction
To enhance scalability and development efficiency, OxyGent introduces a unified atomic abstraction in which every functional element, including agents, tools, and reasoning flows, is encapsulated as a pluggable Oxy node. Figure 1 shows an example of a file management assistant quickly built based on it. This Lego-like approach empowers researchers to rapidly assemble and hot-swap specialized components across diverse scenarios without complex manual reconfiguration. By integrating Aspect-Oriented Programming (AOP) Kiczales et al. (1997), the framework separates core business logic from cross-cutting concerns like logging and security auditing. This design ensures that as the number of agents grows, the system maintains a clean and modular architecture, facilitating seamless horizontal expansion in production settings.
To address the challenge of observability, OxyGent introduces a permission-driven dynamic planning mechanism. Unlike conventional frameworks that rely on static, pre-defined Directed Acyclic Graphs (DAGs), OxyGent defines the potential collaboration space through permission relationships between Oxy nodes, as shown in Figure 1(b). During inference, the system automatically synthesizes the execution path and generates real-time visualizations of the actual call graphs (Figure 1(a)). This transparency allows users to inspect and trace the decision-making trajectories of MAS, ensuring that even in volatile task environments, emergent behaviors remain monitorable and manageable.
To facilitate the agent evolution, OxyGent introduces OxyBank, a specialized platform for AI asset management. OxyBank serves as the evolutionary engine of OxyGent by providing a systematic pipeline that captures online execution traces and converts them into high-quality training samples. By integrating automated rewarding and human-in-the-loop annotation, OxyBank enables the backflow of domain-specific knowledge into the multi-agent system. This closed-loop process supports the joint evolution of agent strategies, ensuring that the collective intelligence of the ecosystem continuously improves through data-centric learning rather than remaining static after deployment. The main contributions are summarized as follows:
  • We propose the unified Oxy abstraction and an AOP-infused orchestration paradigm, enabling Lego-like construction and high scalability for complex multi-agent systems.
  • We implement permission-driven dynamic planning with real-time execution visualization, enabling deep observability into agent behaviors.
  • We introduce OxyBank, a platform for AI asset management that supports agent evolution through automated data backflow and evaluation.
  • Extensive evaluations and case studies demonstrate the effectiveness of OxyGent, proving its robustness in solving collaborative tasks.
@@ 3.3 Evolutionary AI Asset Management via OxyBank
The long-term value of an MAS lies in its ability to self-improve. We introduce OxyBank, an all-in-one AI asset management platform designed to serve as the framework’s memory bank (Figure 6). OxyBank bridges the gap between raw execution data and the continuous self-improvement of the collective agent ecosystem.
The framework supports autonomous evolution by establishing a closed-loop cycle for annotation, auditing, and data backflow. Online execution traces are automatically captured and precipitated as memory assets within OxyBank. These assets are then processed by specialized annotation agents or human experts using customized templates (e.g., QA or business-domain tagging). Once audited, the high-quality samples are used to update knowledge bases or fine-tune models through joint evolution. This pipeline ensures that the MAS continuously refines its decision-making strategies based on validated feedback from the production environment.
@@ 4.2 Case Analysis in Business Scenarios
Beyond academic benchmarks, OxyGent has demonstrated significant industrial impact through wide-scale deployment in enterprise production and open-source communities. To illustrate the framework’s capacity for extreme scale and complexity, we implemented a hierarchical multi-agent classification system for e-commerce customer service, as shown in Figure 8. This system comprises over 2,000 agents tasked with categorizing incoming requests into a taxonomy of more than 2,400 distinct labels. Each category contains as few as 10 samples, making the task a massive few-shot classification problem. The architecture employs a top-down decision chain with dynamic organizational logic and full-chain tracing. Crucially, the system features a topological self-evolution mechanism that autonomously discovers an average of 5.4 new categories per week while retiring obsolete ones. In live production, this MAS-driven approach improved classification accuracy from 61.3% for single-agent-based systems to 85.6%, proving the framework’s robustness in high-concurrency, evolving business environments. Due to space constraints, additional industrial scenarios including automated SOPs, RAG-based assistants, and community-contributed tools are detailed in Appendix A.3.
@@ 5 Conclusion
In this paper, we present OxyGent, an open-source framework for building scalable and self-improving multi-agent systems. By adopting a unified atomic abstraction, OxyGent enables a Lego-like construction process that replaces rigid workflows with permission-driven dynamic planning, offering high observability through real-time execution visualizations. Integrated with OxyBank for AI asset management, the framework establishes a robust pipeline for continuous agent evolution, as evidenced by our extensive evaluations on the GAIA benchmark and industrial applications.

### 29 Apr 2026 v2

@@ Abstract
Deploying production-ready multi-agent systems (MAS) in complex industrial environments remains challenging due to limitations in scalability, observability, and autonomous evolution. We present OxyGent, an open-source framework driven by two core novelties: a unified Oxy abstraction and the OxyBank evolution engine. The unified abstraction encapsulates agents, tools, LLMs, and reasoning flows as pluggable atomic components, enabling Lego-like scalable system composition and non-intrusive monitoring. To enhance observability, OxyGent introduces permission-driven dynamic planning that replaces rigid workflows with execution graphs generated at runtime, providing adaptive visualizations. Furthermore, to support continuous evolution, OxyBank serves as an AI asset management platform that drives automated data backflow, annotation, and joint evolution. Empirical evaluations and real-world case studies show that OxyGent provides a robust and scalable foundation for MAS. OxyGent is fully open-sourced under the Apache License 2.0 at https://github.com/jd-opensource/OxyGent.
@@ 1 Introduction
To enhance scalability and development efficiency, OxyGent introduces a unified abstraction that encapsulates agents, tools, and reasoning flows as pluggable Oxy nodes. This Lego-like approach empowers researchers to assemble and hot-swap specialized components without complex manual reconfiguration rapidly. Furthermore, by integrating Aspect-Oriented Programming (AOP) Kiczales et al. (1997), the framework separates core business logic from cross-cutting concerns, maintaining a clean, modular architecture for seamless expansion in production settings. Building on this abstraction to improve observability, OxyGent replaces static DAGs with permission-driven, dynamic planning. By defining potential collaboration spaces through node permissions (Figure 1(b)), the system automatically synthesizes execution paths at runtime and generates real-time call graph visualizations (Figure 1(a)). This transparency allows users to thoroughly inspect decision-making trajectories, ensuring that emergent MAS behaviors remain fully monitorable and manageable even in volatile task environments.
To facilitate the agent evolution, OxyGent introduces OxyBank, a specialized platform for AI asset management. OxyBank serves as the evolutionary engine of OxyGent by providing a systematic pipeline that captures online execution traces and converts them into high-quality training samples. By integrating automated rewarding and human-in-the-loop annotation, OxyBank enables the backflow of domain-specific knowledge into the multi-agent system. This closed-loop process supports the joint evolution of agent strategies, ensuring that the collective intelligence of the ecosystem continuously improves through data-centric learning rather than remaining static after deployment.
While many recent frameworks emphasize engineering integration, OxyGent’s central research novelties lie in two foundational pillars: the unified Oxy abstraction and the OxyBank evolution engine. The main contributions are summarized as follows:
  • Unified Oxy Abstraction. We propose the unified Oxy abstraction to resolve structural heterogeneity in MAS, enabling Lego-like construction and high scalability.
  • OxyBank Evolution Engine. We introduce an AI asset platform that drives continuous MAS self-improvement via automated data backflow and verifiable evaluation.
  • System Observability. Built upon the Oxy abstraction, we implement permission-driven dynamic planning and an AOP-infused orchestration paradigm, ensuring deep observability.
  • Empirical Validation & Open Source. Evaluations demonstrate OxyGent’s robustness. We release the complete framework, built-in toolkits, a trace visualization Web UI, and comprehensive reference examples.
@@ 3.3 Evolutionary AI Asset Management via OxyBank
The long-term value of an MAS lies in its ability to self-improve. We introduce OxyBank, an AI asset management platform serving as the framework’s memory bank (Figure 6). It bridges the gap between raw execution data and continuous self-improvement through the following mechanisms:
Closed-Loop Asset Pipeline. The framework supports autonomous evolution by establishing a closed-loop cycle for annotation, auditing, and data backflow. Online execution traces are automatically captured and precipitated as memory assets within OxyBank. These assets are processed by annotation agents or experts using customized templates (e.g., QA or business-domain tagging). Once audited, the high-quality samples are used to update knowledge bases or fine-tune models. This pipeline ensures that the MAS continuously refines its decision-making strategies based on validated feedback from production.
Multi-Layered Quality Gating. To address potential degradation from noisy traces, OxyBank implements a strict multi-layered quality-gating pipeline. Specifically, it employs MD5-based deduplication at deposit to prevent frequency bias, and infers trace priorities based on call chain metadata (e.g., end-to-end user interactions are prioritized as P0). Crucially, no raw trace can bypass the strict state machine (pending $\rightarrow$ annotated $\rightarrow$ approved). The pending or rejected data are strictly isolated from the knowledge base, ensuring that only high-fidelity signals drive the evolution process.
Figure 7: AI-driven Optimize Prompt module in OxyGent, which automatically analyzes and refines agent prompts based on historical execution traces, improving prompt quality without manual engineering.
Autonomous Prompt Optimization. Instead of relying solely on manual prompt engineering, OxyGent integrates an AI-driven Optimize Prompt module as shown in Figure 7. It automates the extraction of best practices from validated traces, allowing agents to iteratively refine their system prompts based on historical execution contexts. This dynamically blends the reliability of human-in-the-loop auditing with the scalability of autonomous LLM optimization, significantly accelerating the agents’ adaptation to novel tasks.
@@ 4.1 Evaluations on the GAIA Benchmark
Method Avg Level 1 Level 2 Level 3
Single agent 36.21 61.29 29.56 10.20
+ Multi-agent 42.19 62.37 35.85 24.49
   + Planning 52.16 62.37 54.09 26.53
    + Memory 59.14 77.42 56.60 32.65
Table 1: Ablation study on the GAIA benchmark (%).
Ablation Study. To isolate the performance gains derived from our framework’s specific designs, we conducted an ablation study on the GAIA benchmark (Table 1). Since the unified Oxy abstraction serves as the non-removable foundation of the system, we evaluated the sequential addition of multi-agent orchestration, permission-driven dynamic planning, and the memory backflow (the predecessor to OxyBank). The results show that dynamic planning significantly boosts Level 2 reasoning (+18.24%), while the memory mechanism predominantly enhances L1 task precision (+15.05%) and overall average scores, validating that the gains stem from OxyGent’s architectural designs rather than merely the underlying foundation models.
@@ 4.2 Case Analysis in Business Scenarios
Beyond academic benchmarks, OxyGent has demonstrated significant industrial impact through wide-scale deployment in enterprise production and open-source communities. To illustrate the framework’s capacity for extreme scale and complexity, we implemented a hierarchical multi-agent classification system for e-commerce customer service, as shown in Figure 9.
Setup and Evaluation. This system comprises over 2,000 agents tasked with categorizing incoming requests into a taxonomy of more than 2,400 distinct labels. Each category contains as few as 10 samples, making the task a massive few-shot classification problem. We compared OxyGent against a strong single-agent baseline utilizing RAG retrieval coupled with DeepSeek-R1-Distill-Qwen-32B Guo et al. (2025). The evaluation protocol continuously samples 2,000+ real-world interactions weekly for AI and human joint assessment.
Results and Trade-offs. The MAS-driven approach improved classification accuracy from 61.3% (baseline) to 85.6%. Furthermore, it demonstrated strong topological self-evolution, autonomously discovering and validating an average of 5.4 new categories per week. Regarding cost and latency tradeoffs, the dynamic generation of the 2,000+ agents keeps memory overhead manageable. While the multi-tiered decision architecture increases average inference latency by $2.3\times$ compared to the single-agent baseline, this trade-off is strictly justified for offline business environments prioritizing high precision.
Failure Cases. We observed that highly ambiguous semantic queries occasionally induce repetitive ReAct loops or hallucinations, which are currently mitigated by our trace replay and manual auditing loops. Due to space constraints, additional industrial scenarios including automated SOPs, RAG-based assistants, and community-contributed tools are detailed in Appendix A.3.
5 Discussion
Decision Routing and Structural Resilience. To resolve execution conflicts when multiple nodes qualify, OxyGent utilizes a centralized orchestrator for dynamic, permission-constrained delegation. Under node failures, the framework prevents global trajectory collapse by encapsulating local errors into reasoning memory, enabling autonomous corrective replanning.
Connections to Concurrent Works. AlphaApollo Zhou et al. (2026a) focuses on deep agentic reasoning and learning. In contrast, OxyGent provides a general-purpose, modular, and observable framework, designed for reliable industrial evolution via verifiable data backflow. For inspectability, Landscape of Thoughts Zhou et al. (2026b) provides post-hoc visual analysis of textual reasoning trajectories, whereas OxyGent visualizes multi-agent execution topologically at runtime for immediate architectural debugging. Finally, regarding AR-Bench Zhou et al. (2025), while we evaluated on GAIA to test general multi-modal assistance, OxyGent’s dynamic planning inherently supports the iterative information-seeking required by active reasoning tasks.
@@ 6 Conclusion
In this paper, we present OxyGent, an open-source framework for building scalable and self-improving multi-agent systems driven by two core novelties. First, its unified atomic abstraction enables a Lego-like construction process that replaces rigid workflows with permission-driven dynamic planning, offering high observability through real-time execution visualizations. Second, integrated with OxyBank for AI asset management, the framework continuously refines strategies via a validated feedback pipeline, as evidenced by evaluations on the GAIA benchmark and industrial applications.